# FuzzCoh: Robust Canonical Coherence-Based Fuzzy Clustering of Multivariate Time SeriesLABEL:lastpage

### v1

@@ Abstract
Brain cognitive and sensory functions are often associated with electrophysiological activity at specific frequency bands. Clustering multivariate time series (MTS) data like EEGs is important for understanding brain functions but challenging due to complex non-stationary cross-dependencies, gradual transitions between cognitive states, noisy measurements, and ambiguous cluster boundaries. To address these issues, we develop a robust fuzzy clustering framework in the spectral domain. Our method leverages the Kendall’s tau-based canonical coherence (KenCoh) which extracts meaningful frequency-specific monotonic relationships between groups of channels or regions. KenCoh effectively captures dominant coherence structures while remaining robust against outliers and noise, making it suitable for real EEG datasets that typically contain artifacts. Our method first projects each MTS object onto vectors derived from the KenCoh estimates (i.e, canonical directions), which capture relevant information on the connectivity structure of oscillatory signals in predefined frequency bands. These spectral features are utilized to determine clusters of epochs using a fuzzy partitioning strategy, accommodating gradual transitions and overlapping class structure. Using rigorous numerical experiments, our proposed method is demonstrated to outperform state-of-the-art fuzzy clustering approaches for MTS, including fuzzy C-means and variable-based principal component analysis-based clustering. Lastly, we demonstrate the effectiveness of our approach to EEG data where latent cognitive states such as alertness and drowsiness exhibit frequency-specific dynamics and ambiguity. Our method captures both spectral and spatial features by locating the frequency-dependent structure and brain functional connectivity. Built on the KenCoh framework for fuzzy clustering, it handles the complexity of high-dimensional time series data and is broadly applicable to domains such as neuroscience, wearable sensing, environmental monitoring, and finance.
@@
footnotetext: 1King Abdullah University of Science and Technology (KAUST), Computer, Electrical and Mathematical Sciences and Engineering (CEMSE) Division, Thuwal 23955-6900, Saudi Arabia.
*These authors contributed equally to this work.
Correspondence: ziling.ma@kaust.edu.sa, marasherlin.talento@kaust.edu.sa, ying.sun@kaust.edu.sa, hernando.ombao@kaust.edu.sa
Preprint. Under review.
@@ 1 Introduction
Brain networks, consisting of nodes in brain regions and edges that represent cross-regional interactions or connectivity, are useful to distinguish between control and patient groups, as well as different brain states within an individual (Talento et al., 2024). Clustering brain states based on features extracted from multivariate time series (MTS) of brain signals enables the discovery of distinct neural patterns, identification of cognitive state transitions, and supports personalized, data-driven neuroscience applications (Honey et al., 2007; Van Den Heuvel and Pol, 2010). However, several challenges make this task nontrivial, such as, high dimensionality, temporal dependencies, spectral complexity (Von Luxburg, 2007), and the presence of noise or artifacts (Rodrigues et al., 2017; Ma et al., 2024). These complexity blur the boundaries between brain states and between populations. (Ann Maharaj et al., 2010; Izakian et al., 2015). As a consequence, brain functional networks between brain states within an individual (or between populations, e.g., control and disease) become inseparable.
Aside from the underlying states that switch over time (Samdin et al., 2016; Ting et al., 2017), the multivariate time series is often composed of different oscillatory patterns that reflect different latent conditions in the system. Spectral-domain representations provide a principled way to summarize and analyze such signals. Using spectral analysis, we study the dynamic cross-regional interactions between oscillations which contain meaningful structural and functional information (Van Vugt et al., 2007; Ombao and Pinto, 2024). For instance, electroencephalogram (EEG) oscillations in the 13–30 Hz range (measured in Hertz, or cycles per second) exhibit distinct patterns in patients with Parkinson’s disease (Little and Brown, 2014) and driver’s task-based attention (Li et al., 2022). These findings support the idea that specific frequency bands can offer greater discriminatory power for identifying certain brain states. However, EEG signals, are non-invasive brain imaging modality that are susceptible to artifacts such as eye blinks and external distractions (Jiang et al., 2019). Hence, extracting and categorizing these patterns requires methods that are both robust to noise and flexible in representing overlapping or transitional states.
To address these needs, we propose a robust fuzzy clustering framework in the spectral domain that takes advantage of the robustness of the Kendal’s tau-based canonical coherence (KenCoh) proposed in Talento et al. (2024). KenCoh defined spectral dependence through monotonic canonical relationships between two groups of signals within a specific frequency-band. KenCoh’s ability to quantify spectral dependence across groups of signals enhances the interpretability of multivariate temporal connectivity, e.g., in brain regions where multiple concurrently recorded neural signals reflect localized brain functions. In other words, this technique enables the extraction of coherence-based features that are both interpretable and resistant to distortion by artifacts or outliers.
Figure 1: Schematic diagram of the FuzzCoh method. Electroencephalogram (EEG) signals collected from multiple nodes (left panel) are processed through the FuzzCoh pipeline to extract meaningful connectivity patterns.
In our proposed method, spectral features extracted from KenCoh are integrated into a fuzzy clustering framework—hence the name FuzzCoh. Our method permits partial membership assignments that are frequency-based, thereby effectively capturing transitional or mixed-state phenomena frequently observed in real-world datasets. In this article, we demonstrate the effectiveness of the proposed FuzzCoh method using the multi-node, multi-trial EEG data designed to detect whether an individual is in a drowsy or alert state (Cao et al., 2019). Detecting drowsiness is an especially relevant application, as it has significant safety implications, that is, drowsy driving is implicated in a large fraction of road accidents and fatalities worldwide (Tefft, 2010; Higgins et al., 2017; Filtness et al., 2017). Drowsiness and alertness are not binary brain states. In fact, there is “fuzziness” between drowsiness and alertness. Drivers frequently experience intermediate or transitional phases of cognitive fatigue (Albadawi et al., 2022). This ambiguity of states makes fuzzy clustering a well-suited approach for modeling the underlying dynamics.
Figure 1 shows the placement of 14 EEG channels on the scalp, including regions near the prefrontal lobe (FP1, FP2), frontal lobe (F3, F4, F7, F8), temporo-parietal lobe (TP7, TP8, T7, T8, P3, P4), and occipital lobe (O1, O2). These regions are associated with brain functionalities that are relevant to driving-task Li et al. (2022). EEG is widely used for monitoring cognitive states, and its spectral properties, particularly in the Delta $\Omega=(0,4]$ Hz, Theta $\Omega=(4,8]$ Hz, Alpha $\Omega=(8,12]$ Hz, Beta $\Omega=(12,30]$ Hz, and Gamma $\Omega=(30,50]$ Hz bands, which are closely linked to mental alertness and drowsiness (Klimesch, 1999; Correa et al., 2014; Chaabene et al., 2021). Here, these frequency bands (e.g., Delta-band) are extracted from the signal or time series through filtering (i.e., Butterworth filter) (Ombao and Pinto, 2024). Let $\{X(t)\}_{t=1}^{T}$ be a weakly-stationary time-series, and let $g(h)$, for $h\in\mathbb{Z}$, be an absolutely summable linear filter, such that, $\sum_{h=-\infty}^{\infty}|g(h)|<\infty$. Then, the filtered series is denoted as $X_{\Omega}(t):=\sum_{h=-\infty}^{\infty}g(t-h)X(t)$. This filtered series $X_{\Omega}(t)$ has zero spectral power outside $\Omega$-band (see Ombao and Pinto, 2024, for details).
While EEG serves as our main case study, the proposed framework is broadly applicable to any domain where frequency-specific dependencies and ambiguous class boundaries are present. Examples include wearable sensor streams, financial volatility cycles, or environmental time series.
The main contributions of this work are as follows:
  1. 1.
    We develop a novel fuzzy clustering method in the frequency domain, called FuzzCoh, which identifies frequency bands (or underlying oscillations) that most effectively discriminate between MTS states based on their dependence structure.
  2. 2.
    We adopt a robust estimation of multivariate spectral-dependence structure, enabling our method to remain stable and accurately represent the underlying structure of the data, even under severe contamination.
  3. 3.
    We demonstrate the advantages of our method on EEG data, where transitional brain states are present and rigid group assignments from hard clustering can be misleading. Additionally, our method provides a pathway for relating brain states to the underlying brain functional-connectivity structure.
  4. 4.
    Our method incorporates spatial information, recognizing that the brain consists of distinct regions. By examining the connectivity between regions, we can identify which brain areas show the most pronounced differences between alert and drowsy states.
In summary, our novel rank-based fuzzy clustering approach is broadly applicable to real-world MTS problems, where spatial and spectral information is informative and non-ignorable overlaps in the underlying states are present.
@@ 2.1 Lagged dependence matrix of filtered signals
Consider a non-stationary MTS $\boldsymbol{Z}_{BT\times(p+q)}=(\boldsymbol{X}_{BT\times p},\boldsymbol{Y}_{BT\times q})$. Suppose that this process is ‘locally stationary’ so that for a block $b,(b=1,\dots,B)$ of length $T$, the time series are approximately stationary. Let $\boldsymbol{Z}^{(b)}_{\Omega}(t)\in\mathbb{R}^{p+q}$ denote the $\boldsymbol{Z}_{t}$ corresponding to the $b$-th block filtered to the $\Omega$ frequency band. We define $\boldsymbol{P}^{(b)}(\Omega,\ell)$ as the $b$-th block lagged dependence among filtered signals on $\Omega$ at lag $\ell=0,\pm 1,\dots,\pm L$, i.e., ${P}^{(b)}_{jk}(\Omega,\ell)$ = $f(Z_{\Omega,j}(t),Z_{\Omega,k}(t+\ell))$ = $f(Z_{\Omega,j}(t-\ell),Z_{\Omega,k}(t))$, for $j,k=1,\dots,p+q$ (see Talento et al., 2024, for details). The function $f:\mathbb{R}^{2}\rightarrow(-1,1)$ is defined for $t\neq s=1,\dots,T-\ell$ as
${\tau}^{(b)}(\Omega,\ell)=\mathbb{E}[\text{sign}\{(Z^{(b)}_{\Omega}(t)-Z^{(b)}_{\Omega}(s))(Z^{(b)}_{\Omega}(t-\ell)-Z^{(b)}_{\Omega}(s-\ell))\}],$
$f(Z_{\Omega,j}(t-\ell),Z_{\Omega,k}(t))=\sin\left(\frac{\pi}{2}{\tau}^{(b)}(\Omega,\ell)\right).$ (1)
We can subset the lagged dependence matrix into four sub-parts as
$\boldsymbol{P}^{(b)}(\Omega,\ell)=\begin{pmatrix}\boldsymbol{P}^{(b)}_{X,X}(\Omega,\ell)&\boldsymbol{P}^{(b)}_{X,Y}(\Omega,\ell)\\ \boldsymbol{P}^{(b)}_{Y,X}(\Omega,\ell)&\boldsymbol{P}^{(b)}_{Y,Y}(\Omega,\ell)\end{pmatrix}.$ (2)
These four submatrices of $\boldsymbol{P}^{(b)}(\Omega,\ell)$ are the ingredients in identifying fuzzy series. We now define spectral canonical coherence at block $b$, denoted as $R_{1}(b,\Omega)$. The overall measure of spectral association between $\boldsymbol{X}(t)$ and $\boldsymbol{Y}(t)$ is defined as
$R_{1}(b,\Omega)=\max_{\boldsymbol{u}_{b,\Omega},\boldsymbol{v}_{b,\Omega},l}\left\{\boldsymbol{u}^{\top}_{b,\Omega}\boldsymbol{P}^{(b)}_{X,Y}(\Omega,\ell)\boldsymbol{v}_{b,\Omega}\right\}^{2},\;\;\;\text{such that,}$, (3)
$\boldsymbol{u}^{\top}_{b,\Omega}\boldsymbol{P}^{(b)}_{X,X}(\Omega,0)\boldsymbol{u}_{b,\Omega}=\boldsymbol{v}^{\top}_{b,\Omega}\boldsymbol{P}^{(b)}_{Y,Y}(\Omega,0)\boldsymbol{v}_{b,\Omega}=1.$
These $\boldsymbol{u}_{b,\Omega}$ and $\boldsymbol{v}_{b,\Omega}$ are the block-wise canonical variates (see Talento et al., 2024, for details). These vectors show the relative contribution of the variables to the maximum association, $R_{1}(b,\Omega)$. This paper exploited the structure of connectivity through $\boldsymbol{u}_{b,\Omega}$ and $\boldsymbol{v}_{b,\Omega}$ to determine the clusters of multiple time series for a specific oscillation band.
2.2 FuzzCoh: The robust fuzzy clustering algorithm based on KenCoh
@@ 2.2.1 The clustering procedures
We use the standardized canonical directions, $\boldsymbol{u}_{b,\Omega}$ and $\boldsymbol{v}_{b,\Omega}$, which are computed from the filtered signals $\boldsymbol{X}^{(b)}_{\Omega}\in\mathbb{R}^{T\times p}$ and $\boldsymbol{Y}^{(b)}_{\Omega}\in\mathbb{R}^{T\times q}$, and capture the maximum cross-dependence at a given spectral band $\Omega$, as spectral features to conduct the fuzzy clustering.
Consider a filtered MTS dataset consisting of $B$ realizations, denoted by $\boldsymbol{Z}_{\Omega}=\{\boldsymbol{Z}^{(1)}_{\Omega},\dots,\boldsymbol{Z}^{(B)}_{\Omega}\}$. Also, let $\boldsymbol{Z}^{(b)}_{\Omega}$ be the concatenation of realizations $\boldsymbol{X}^{(b)}_{\Omega}\in\mathbb{R}^{T\times p}$ and $\boldsymbol{Y}^{(b)}_{\Omega}\in\mathbb{R}^{T\times q}$, that is,
$\boldsymbol{Z}^{(b)}_{\Omega}=\left(\boldsymbol{X}^{(b)}_{\Omega},\boldsymbol{Y}^{(b)}_{\Omega}\right),$
[…]
$&\min_{\overline{\boldsymbol{d}}^{\Omega},\boldsymbol{E}}\sum_{b=1}^{B}\sum_{c=1}^{C}(e_{bc}^{\Omega})^{m}\left\|\boldsymbol{d}_{b}^{\Omega}-\overline{\boldsymbol{d}}_{c}^{\Omega}\right\|^{2},\\ &\text{subject to}\quad\sum_{c=1}^{C}e_{bc}^{\Omega}=1,\quad e_{bc}^{\Omega}\geq 0,\quad\forall b=1,\dots,B.\end{aligned}\right.$ (4)
@@
$e_{bc}^{\Omega}=\left(\sum_{c^{\prime}=1}^{C}\left(\frac{\left\|\boldsymbol{d}_{b}^{\Omega}-\overline{\boldsymbol{d}}_{(c)}^{\Omega}\right\|^{2}}{\left\|\boldsymbol{d}_{b}^{\Omega}-\overline{\boldsymbol{d}}_{(c^{\prime})}^{\Omega}\right\|^{2}}\right)^{\frac{1}{m-1}}\right)^{-1},$ (5)
@@
$\overline{\boldsymbol{d}}_{c}^{\Omega}=\frac{\sum_{b=1}^{B}(e_{bc}^{\Omega})^{m}\boldsymbol{d}_{b}^{\Omega}}{\sum_{b=1}^{B}(e_{bc}^{\Omega})^{m}}.$ (6)
2.2.2 The cluster validity index for the automatic selection of $C$ and $m$
As stated by Rhee and Oh (1996), if fuzzy cluster analysis is to make a significant contribution to engineering applications, greater attention must be given to the fundamental decision of determining the number of clusters in the data. In this paper, we present a cluster validity index (CVI) to help determine the optimal selection of $C$. We adopt the Fuzzy Silhouette index (FSI) (Rawashdeh and Ralescu, 2012), which generalizes the classical silhouette index (Rousseeuw, 1987) to the fuzzy setting. For each MTS object $d_{b}^{\Omega}$, the silhouette value with respect to cluster $c$ is computed as
$s_{bc}^{\Omega}=\frac{n^{\Omega}_{bc}-a^{\Omega}_{bc}}{\max(a^{\Omega}_{bc},n^{\Omega}_{bc})},$ (7)
where $a_{bc}$ is the average dissimilarity between $d_{b}^{\Omega}$ and all other objects in cluster $c$, weighted by their membership degrees, defined as,
$a^{\Omega}_{bc}=\frac{\sum_{j\neq b}(e_{jc}^{\Omega})^{m}\cdot\lVert d_{b}^{\Omega}-d_{j}^{\Omega}\rVert}{\sum_{j\neq b}(e_{jc}^{\Omega})^{m}},$ (8)
and $n_{bc}$ is the minimum average dissimilarity between $d_{b}^{\Omega}$ and all objects in another cluster $c^{\prime}\neq c$, defined as,
$n^{\Omega}_{bc}=\min_{c^{\prime}\neq c}\left(\frac{\sum_{b=1}^{B}(e_{bc^{\prime}}^{\Omega})^{m}\cdot\lVert d_{b}^{\Omega}-d_{j}^{\Omega}\rVert}{\sum_{b=1}^{B}(e_{bc^{\prime}}^{\Omega})^{m}}\right).$ (9)
The overall FSI, for a specific frequency band $\Omega$, is then given by,

### v2

# Robust Spectral Fuzzy Clustering of Multivariate Time Series with Applications to ElectroencephalogramLABEL:lastpage

@@ Abstract
Clustering multivariate time series (MTS) is challenging due to non-stationary cross-dependencies, noise contamination, and gradual or overlapping state boundaries. We introduce a robust fuzzy clustering framework in the spectral domain that leverages Kendall’s tau–based canonical coherence to extract frequency-specific monotonic relationships across variables. Our method takes advantage of dominant frequency-based cross-regional connectivity patterns to improve clustering accuracy while remaining resilient to outliers, making the approach broadly applicable to noisy, high-dimensional MTS. Each series is projected onto vectors generated from a spectral matrix specifically tailored to capture the underlying fuzzy partitions. Numerical experiments demonstrate the superiority of our framework over existing methods. As a flagship application, we analyze electroencephalogram recordings, where our approach uncovers frequency- and connectivity-specific markers of latent cognitive states such as alertness and drowsiness, revealing discriminative patterns and ambiguous transitions.
@@ 1 Introduction
Brain networks, composed of nodes representing brain regions and edges representing cross-regional connectivity, are central to understanding neural dynamics and distinguishing between cognitive states and populations (Wu et al., 2025). Clustering multivariate time series (MTS) of brain signals enables the discovery of latent neural patterns, detection of cognitive transitions, and development of personalized neuroscience applications (Honey et al., 2007; Van Den Heuvel and Pol, 2010). Electroencephalogram (EEG) signals are particularly challenging: they are high-dimensional, temporally dependent, spectrally complex (Von Luxburg, 2007), and contaminated by artifacts such as eye blinks or muscle activity (Rodrigues et al., 2017; Jiang et al., 2019). Moreover, brain states (such as alertness and drowsiness) often evolve gradually with overlapping boundaries (Masulli et al., 2019; Zhao et al., 2024), making fuzzy clustering more appropriate than crisp partitioning.
Spectral-domain analysis provides a principled way to summarize oscillatory interactions and capture functional connectivity across regions (Van Vugt et al., 2007; Ombao and Pinto, 2024). EEG oscillations in specific bands have been linked to neurological disorders (Little and Brown, 2014) and task-related attention (Li et al., 2022), underscoring the discriminative power of frequency-specific features. Thus, robust fuzzy clustering methods that incorporate spectral and spatial information are highly desirable for EEG analysis.
However, existing fuzzy clustering approaches for MTS remain limited. Most operate in the time domain, relying on temporal features, autoregressive models, or correlation structures (Izakian et al., 2015; Ann Maharaj et al., 2010). Robustness has been considered through trimming strategies (D’Urso et al., 2023) or quantile cross-spectral features (López-Oriona et al., 2022), but these are computationally intensive. Subspace-based methods such as Ma et al. (2025) improve scalability but still neglect cross-regional connectivity. Overall, current methods are either time-domain only, sensitive to contamination, or fail to leverage spatial connectivity—leaving a gap for a robust, frequency-domain fuzzy clustering framework that can handle overlapping EEG states.
We propose FuzzCoh, a fuzzy clustering framework in the spectral domain that leverages Kendall’s tau–based canonical coherence (KenCoh) (Talento et al., 2025) to extract robust connectivity features. FuzzCoh allows partial membership assignments, capturing transitional or mixed-state phenomena, and provides interpretable insights into frequency-specific and connectivity-specific dependencies. We validate FuzzCoh on multi-node, multi-trial EEG data for detecting drowsy versus alert states (Cao et al., 2019), a problem with significant safety implications (Tefft, 2010; Higgins et al., 2017; Filtness et al., 2017). Since drowsiness and alertness are not strictly binary but instead exhibit fuzzy transitional phases (Albadawi et al., 2022; Ma et al., 2025), fuzzy clustering is particularly well suited. [^1] To demonstrate that FuzzCoh is not limited to EEG data, we provide an additional application on motion MTS data in the Appendix, Section 6.1. Our contributions are threefold: (1) we develop a novel frequency-domain fuzzy clustering method, FuzzCoh, which identifies oscillatory bands that discriminate between different MTS states; (2) we incorporate robust estimation of spectral dependence to maintain stability under contamination; and (3) by examining cross-regional connectivity, we account for spatial information and link brain states to functional networks, thereby providing deeper neuroscientific insights.
@@ 2.1 Lagged dependence matrix of filtered signals
Let $\bm{Z}_{BT\times(p+q)}=(\bm{X}_{BT\times p},\bm{Y}_{BT\times q})$} be a non-stationary MTS. Suppose that this process is ‘locally stationary’ so that for a block $b,(b=1,\dots,B)$ of length $T$, the time series are approximately stationary. Denote $\bm{Z}^{(b)}_{\Omega}(t)\in\mathbb{R}^{p+q}$} as the $\bm{Z}_{t}$ corresponding to the $b$-th block filtered to the $\Omega$ frequency band. Furthermore, let
$\zeta_{0,\Omega}^{(b)}=Z^{(b)}_{\Omega}(t)-Z^{(b)}_{\Omega}(s),$
$\zeta_{\ell,\Omega}^{(b)}=Z^{(b)}_{\Omega}(t-\ell)-Z^{(b)}_{\Omega}(s-\ell),$
${\tau}^{(b)}(\Omega,\ell)=\mathbb{E}[\text{sign}\{\zeta_{0,\Omega}^{(b)}\zeta_{\ell,\Omega}^{(b)}\}],$
$f(Z_{\Omega,j}(t-\ell),Z_{\Omega,k}(t)):=\sin\left(\frac{\pi}{2}{\tau}^{(b)}(\Omega,\ell)\right).$ (1)
for $t\neq s$ and a fixed $\ell$, i.e., $f:\mathbb{R}^{2}\rightarrow(-1,1)$. Talento et al. (2025) define the $b$-th block lagged dependence matrix among filtered signals on $\Omega$ at lag $\ell=0,\pm 1,\dots,\pm L$} as
$\bm{P}^{(b)}(\Omega,\ell)=\begin{pmatrix}\bm{P}^{(b)}_{X,X}(\Omega,\ell)&\bm{P}^{(b)}_{X,Y}(\Omega,\ell)\\ \bm{P}^{(b)}_{Y,X}(\Omega,\ell)&\bm{P}^{(b)}_{Y,Y}(\Omega,\ell)\end{pmatrix}$, (2)
$\text{where}\ {P}^{(b)}_{jk}(\Omega,\ell)$ $:=f(Z_{\Omega,j}(t),Z_{\Omega,k}(t+\ell))$
$=f(Z_{\Omega,j}(t-\ell),Z_{\Omega,k}(t))$
for $j,k=1,\dots,p+q$. We then extract the cross-vector topologies at block $b$ and frequency-band $\Omega$, denoted as $\bm{u}_{b,\Omega}$ and $\bm{v}_{b,\Omega}$, through the following objective function in Talento et al. (2025), and denoted as $g^{\Omega}(b)$, i.e.,
$g^{\Omega}(b):=\max_{\bm{u}_{b,\Omega},\bm{v}_{b,\Omega},\ell}\left\{\bm{u}^{\top}_{b,\Omega}\bm{P}^{(b)}_{X,Y}(\Omega,\ell)\bm{v}_{b,\Omega}\right\}^{2},\ \text{such that}$
$\bm{u}^{\top}_{b,\Omega}\bm{P}^{(b)}_{X,X}(\Omega,0)\bm{u}_{b,\Omega}=\bm{v}^{\top}_{b,\Omega}\bm{P}^{(b)}_{Y,Y}(\Omega,0)\bm{v}_{b,\Omega}=1.$
These $\bm{u}_{b,\Omega}$ and $\bm{v}_{b,\Omega}$, called the block-wise canonical variates (see Talento et al., 2025, for details), show the contribution of the variables relative to $g^{\Omega}(b)$. Thus, this paper exploited the connectivity-structure through $\bm{u}_{b,\Omega}$ and $\bm{v}_{b,\Omega}$, and determine the clusters of multiple time series for a specific oscillation band.
2.2 FuzzCoh: The algorithm
@@ 2.2.1 The clustering procedures
We use the standardized canonical directions, $\bm{u}_{b,\Omega}$ and $\bm{v}_{b,\Omega}$, which are computed from the filtered signals $\bm{X}^{(b)}_{\Omega}\in\mathbb{R}^{T\times p}$ and $\bm{Y}^{(b)}_{\Omega}\in\mathbb{R}^{T\times q}$, and capture the maximum cross-dependence at a given spectral band $\Omega$, as spectral features to conduct the fuzzy clustering.
Consider a filtered MTS dataset consisting of $B$ realizations, denoted by $\bm{Z}_{\Omega}=\{\bm{Z}^{(1)}_{\Omega},\dots,\bm{Z}^{(B)}_{\Omega}\}$, where
$\bm{Z}^{(b)}_{\Omega}=\left(\bm{X}^{(b)}_{\Omega},\bm{Y}^{(b)}_{\Omega}\right)$,
for $b=1,\dots,B$. Denote $\bm{d}_{b}^{\Omega}=(\lvert\bm{u}_{b,\Omega}^{\top}\rvert,\lvert\bm{v}_{b,\Omega}^{\top}\rvert)^{\top}\in\mathbb{R}^{p+q},$} where $\lvert\cdot\rvert$ is the absolute-value operator, applied element-wise to a vector. It concatenates the element-wise absolute values of the standardized canonical directions obtained at frequency band $\Omega$ and $b$-th series. We compute this for each MTS object to obtain a collection of $\{\bm{d}_{b}^{\Omega}\}_{b=1}^{B}$, denoted as $\bm{d}^{\Omega}=\{\bm{d}_{1}^{\Omega},\cdots,\bm{d}_{B}^{\Omega}\}$. Then we perform a fuzzy $C$-means clustering model using $\bm{d}^{\Omega}$, aiming to find a set of cluster centers $\overline{\bm{d}}^{\Omega}=\left\{\overline{\bm{d}}_{1}^{\Omega},\dots,\overline{\bm{d}}_{C}^{\Omega}\right\}$, and the $B\times C$ fuzzy coefficient matrix, $\bm{E}=(e_{bc}^{\Omega})$, for $b=\{1,\cdots,B\}$} and $c=\{1,\cdots,C\}$, by solving the minimization problem
$\left\{\begin{aligned} &\min_{\overline{\bm{d}}^{\Omega},\bm{E}}\sum_{b=1}^{B}\sum_{c=1}^{C}(e_{bc}^{\Omega})^{m}\left\|\bm{d}_{b}^{\Omega}-\overline{\bm{d}}_{c}^{\Omega}\right\|^{2},\\ &\text{subject to}\quad\sum_{c=1}^{C}e_{bc}^{\Omega}=1,\quad e_{bc}^{\Omega}\geq 0,\quad\forall b=1,\dots,B.\end{aligned}\right$. (3)
@@
$e_{bc}^{\Omega}=\left(\sum_{c^{\prime}=1}^{C}\left(\frac{\left\|\bm{d}_{b}^{\Omega}-\overline{\bm{d}}_{(c)}^{\Omega}\right\|^{2}}{\left\|\bm{d}_{b}^{\Omega}-\overline{\bm{d}}_{(c^{\prime})}^{\Omega}\right\|^{2}}\right)^{\frac{1}{m-1}}\right)^{-1},$ (4)
@@
$\overline{\bm{d}}_{c}^{\Omega}=\frac{\sum_{b=1}^{B}(e_{bc}^{\Omega})^{m}\bm{d}_{b}^{\Omega}}{\sum_{b=1}^{B}(e_{bc}^{\Omega})^{m}}.$ (5)
2.2.2 Cluster validity index for $C$ and $m$
As stated by Rhee and Oh (1996), if fuzzy cluster analysis is to make a significant contribution to engineering applications, greater attention must be given to the fundamental decision of determining the number of clusters in the data. In this paper, we present a cluster validity index to help determine the optimal selection of $C$ and $m$. We adopt the Fuzzy Silhouette index (FSI) (Rawashdeh and Ralescu, 2012), which generalizes the classical silhouette index (Rousseeuw, 1987) to the fuzzy setting. For each MTS object $d_{b}^{\Omega}$, the silhouette value with respect to cluster $c$ is computed as
$s_{bc}^{\Omega}=\frac{n^{\Omega}_{bc}-a^{\Omega}_{bc}}{\max\left(a^{\Omega}_{bc},n^{\Omega}_{bc}\right)},$ (6)
where $a^{\Omega}_{bc}$ is the average dissimilarity between $\bm{d}_{b}^{\Omega}$ and all other objects in cluster $c$, weighted by their membership degrees, defined as
$a^{\Omega}_{bc}=\frac{\sum_{\begin{subarray}{c}j=1\\ j\neq b\end{subarray}}^{B}(e_{jc}^{\Omega})^{m}\cdot\lVert\bm{d}_{b}^{\Omega}-\bm{d}_{j}^{\Omega}\rVert}{\sum_{\begin{subarray}{c}j=1\\ j\neq b\end{subarray}}^{B}(e_{jc}^{\Omega})^{m}},$ (7)
and $n^{\Omega}_{bc}$ is the minimum average dissimilarity between $\bm{d}_{b}^{\Omega}$ and all objects in another cluster $c^{\prime}\neq c$, defined as
$n^{\Omega}_{bc}=\min_{c^{\prime}\neq c}\left(\frac{\sum_{\begin{subarray}{c}j=1\\ j\neq b\end{subarray}}^{B}(e_{jc^{\prime}}^{\Omega})^{m}\cdot\lVert\bm{d}_{b}^{\Omega}-\bm{d}_{j}^{\Omega}\rVert}{\sum_{\begin{subarray}{c}j=1\\ j\neq b\end{subarray}}^{B}(e_{jc^{\prime}}^{\Omega})^{m}}\right).$ (8)
The overall FSI, for a specific frequency band $\Omega$, is then given by